A mercury thermometer is constructed as shown above. The glass capillary tube has a diameter of 0.00400 cm, and the bulb has a diameter of 0.250 cm. Find the change in height of the mercury column that occurs with a temperature change of 20oC. Linear expansion coefficient of glass is 4 x 10-6 ( oC)-1 Volumetric expansion coefficient of mercury is 180 x 10-6 ( oC)-1 HINT: Assume that the dimensions of the capillary tube will not change (only the bulb experiences temperature change)


1
Expert's answer
2021-09-28T16:55:23-0400
"\pi r_t^2\Delta h=\frac{4}{3}\pi r_b^3\beta \Delta T\\
\pi (0.002)^2\Delta h=\frac{4}{3}\pi (0.175)^3(180\cdot10^{-6})(30)\\\Delta h=9.65\ cm"